# WiFo-M2: Plug-and-Play Multi-Modal Sensing via Foundation Model to Empower Wireless Communications

### v1

@@ IV-A2 LiDAR Processing Pipeline
IV-B Network Architecture and Working Flow of WiFo-M2
@@
IV-B1 CSI Feature Extraction
For CSI feature extraction, we employ the pre-trained WiFo model [21] as a powerful spatio-temporal-frequency (STF) feature extractor. Given an input CSI data $\mathbf{H}\in\mathbb{C}^{K\times N}$, the WiFo model processes it through its transformer-based encoder. The output is a fixed-dimensional feature vector $\mathbf{z}^{\text{C}}\in\mathbb{R}^{d}$.
IV-B2 Image and LiDAR Feature Extraction
@@
where $f_{\text{ResNet}}$ and $f_{\text{PointNet}}$ denote the ResNet and PointNet backbone networks respectively. We denote the frame-level features for image and LiDAR within a time window $T-n_{\rm I}\Delta t_{\rm I},T]$ or $T-n_{\rm L}\Delta t_{\rm L},T]$ as $\bm{u}_{t}^{\text{I}}\in\mathbb{R}^{d}$ and $\bm{u}_{t}^{\text{L}}\in\mathbb{R}^{d}$, respectively.
Stage 2: Temporal Feature Modeling. To capture temporal dependencies, both modalities employ the GRU network that learns the underlying dynamics of the environment. The GRU processes the sequences of intermediate frame-level features $\{\bm{u}^{\rm x}_{t}\}_{t=T-n_{\rm x}\Delta t_{\rm x}}^{T},{\rm x}\in\{\rm L,\rm I\}$ to produce a sequence of hidden states, which are then projected to the final feature space. Let $\Delta\tau$ be the CSI estimation interval. The resulting projected feature sequence corresponds to the prediction window $T,T+\Delta\tau,\cdots,T+\Delta t_{\rm x}]$, i.e., from the current sensing timestamp $T$ up to the moment when the next sensing frame is captured. This operation can be expressed as:
@@
Stage 3: Multi-Timestamp Feature Generation and Alignment. A key design of WiFo-M2’s architecture is that the GRU generates features for multiple future timestamps simultaneously. During pre-training, the sequence of final features $\{\bm{v}^{\rm x}_{\tau}\}_{\tau=T}^{T+\Delta t_{\rm x}},{\rm x}\in\{\rm L,\rm I\}$ generated by WiFo-M2 are individually aligned with the CSI feature at the corresponding timestamp. For a given timestamp $\tau$, a certain sensing feature $\bm{v}^{\rm{x}}_{\tau}$ is selected and paired with the CSI feature $\bm{z}^{\text{C}}_{\tau}$ from the same timestamp to form a positive pair for CL (detailed in the next section). As shown in Fig. 4, this design enables WiFo-M2 to learn to generate temporally aligned features at specific future timestamps, which can then be directly compared with CSI features from the corresponding time in the CL objective, ensuring precise cross-modal temporal alignment.
IV-C Pre-training Strategy for WiFo-M2: ContraSoM
This section presents the proposed ContraSoM pre-training strategy in detail, covering the modality-specific data-augmentation techniques together with the overall contrastive learning pipeline. By aligning multi-modal sensing and wireless channel in the latent space, ContraSoM essentially performs the SoM processing outlined in [2].
IV-C1 Modality-Specific Data Augmentation
@@
The augmentation process begins with a forward noising step. Taking the final LiDAR feature at a specific time step $i$, denoted as $\bm{v}^{\rm L}_{i}$, as an example, we gradually inject Gaussian noise into this original feature over a series of $M$ discrete steps. This forward noising process is defined as a Markov chain:
@@
$\bm{v}^{\rm L}_{m}=\sqrt{\bar{\alpha}_{m}}\,\bm{v}^{\rm L}_{i}+\sqrt{1-\bar{\alpha}_{m}}\,\epsilon,\quad\epsilon\sim\mathcal{N}(0,\mathbf{I})$ (11)
@@
For efficient feature generation, we employ the Denoising Diffusion Implicit Models (DDIM) sampling during the reverse process, where the noise predictor $\epsilon_{\theta}(\cdot)$ is a learnable NN. Starting from a randomly sampled noisy feature $\bm{v}^{\rm L}_{m}$, the DDIM sampler with stride $\Delta m=\max(\lfloor m/u\rfloor,1)$ (where $u=8$) efficiently recovers a cleaned feature in fewer steps:
$\bm{v}^{\rm L}_{m-\Delta m}=$ $\sqrt{\bar{\alpha}_{m-\Delta m}}\left(\frac{\bm{v}^{\rm L}_{m}-\sqrt{1-\bar{\alpha}_{m}}\,\epsilon_{\theta}(\bm{v}^{\rm L}_{m},m)}{\sqrt{\bar{\alpha}_{m}}}\right)$
$+\sqrt{1-\bar{\alpha}_{m-\Delta m}}\,\epsilon_{\theta}(\bm{v}^{\rm L}_{m},m).$ (12)
After $U=24$ reverse steps, we obtain the augmented feature $\tilde{\bm{v}}^{\rm L}_{i}$. To enhance diversity, we generate two distinct augmented views $\tilde{\bm{v}}_{i}^{\rm L(1)}$ and $\tilde{\bm{v}}_{i}^{\rm L(2)}$ by starting from different noise samples:
$\tilde{\bm{v}}_{i}^{\rm L(1)}=\text{DDIM}(\bm{v}^{\rm L}_{i},\epsilon_{1},m_{1}),\quad\tilde{\bm{v}}_{i}^{\rm L(2)}=\text{DDIM}(\bm{v}^{\rm L}_{i},\epsilon_{2},m_{2}),$
where $m_{1},m_{2}\sim\mathcal{U}\{200,400\}$ represent the initial diffusion steps, and $\epsilon_{1},\epsilon_{2}\sim\mathcal{N}(0,\mathbf{I})$ are two independent Gaussian noise vectors. The two independent pairs $(m_{1},\epsilon_{1})$ and $(m_{2},\epsilon_{2})$ produce two distinct noised versions of the original feature, which act as separate deterministic starting points for the DDIM reverse process and ultimately yield the two complementary augmented views.
IV-C2 Contrastive Learning Objectives in WiFo-M2
@@
For the LiDAR modality, we incorporate both the original LiDAR feature $\bm{v}^{\rm L}$ and the two diffusion-augmented features $\tilde{\bm{v}}^{\rm L(1)}$, $\tilde{\mathbf{v}}^{\rm L(2)}$ into the contrastive learning framework. The overall objective consists of two components: a contrastive loss $\mathcal{L}^{\rm L}_{\text{contra}}$ and a diffusion training loss $\mathcal{L}_{\text{diff}}$. Specifically, we formulate a triplet contrastive loss to enforce consistency between three pairs: the original LiDAR features and CSI features, and the augmented LiDAR features and CSI features:
@@
$\mathcal{L}_{\text{diff}}=\mathbb{E}_{t,\mathbf{z}_{0},\epsilon}\left[\|\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},t)\|^{2}\right],$ (17)
with $t\sim\mathcal{U}\{1,\dots,T\}$ and $\epsilon\sim\mathcal{N}(0,\mathbf{I})$.
@@ V-A Datasets
Pre-training Dataset M3SC [27] Intersection Traffic $28$ I1 $32\times 1$ $0.2$ $512$ $881$
I2 $64\times 1$ 11 $1024$ $868$
I3 $16\times 8$ $0.8$ $512$ $886$
I4 $16\times 8$ $0.4$ $512$ $896$
I5 $8\times 8$ $0.8$ $1024$ $878$
I6 $8\times 4$ 11 $256$ $871$
SynthSoM [28] Dense Building $4.95$ B1 $16\times 16$ $0.05$ $256$ $2270$
B2 $16\times 8$ $0.05$ $512$ $2098$
B3 $8\times 8$ $0.05$ $512$ $1852$
B4 $8\times 4$ $0.1$ $512$ $2353$
B5 $4\times 4$ $0.1$ $1024$ $2379$
@@
We construct a multi-modal pre-training dataset that includes LiDAR, RGB images, and CSI data. The CSI data is collected from the links between the BS and the users. In all datasets, the users are equipped with a single antenna, while the BS employs either a UPA or ULA configuration, with adjacent antenna spacing set to half the wavelength at the central frequency. The data is sourced from the dense building scenario in the SynthSoM dataset [28] and the intersection traffic scenario in the M3SC dataset [27]. Note that various communication system configurations are considered, covering a wide range of subcarrier numbers, bandwidths, and antenna sizes. The detailed simulation configurations of the pre-training datasets are presented in Table I.
We also construct a testing set, with detailed simulation configurations shown in Table II. The validation set consists of two parts: $7$ test datasets for seen scenarios and $3$ datasets for unseen scenarios. First, to validate the in-distribution performance of WiFo-M2, we construct datasets L1 to L7, which are selected from the same links as the pre-training datasets. Among these, datasets L1, L5, and L6 are formed by merging several links with the same number of antennas and subcarriers but smaller data volumes. To test the cross-scenario generalization ability of WiFo-M2, we construct datasets C1, S1, and V1 based on three new scenarios not included in the pre-training datasets, namely, the campus scenario in the SynthSoM-Twin dataset [30], the dist_cam scenario in the ViWi dataset [31], and the Scenario 31 in the DeepSense-6G dataset [32], respectively. Notably, these three datasets operate at entirely new carrier frequencies, and their images and LiDAR point clouds exhibit visual appearances and spatial layouts unseen in the pre-training data. This pronounced gap between training and evaluation domains highlights the strong universality and generalization capability of WiFo‑M2. Fig. 5 provides an overview of the multi-modal data examples from all the aforementioned scenarios.
@@ V-B Network and Pre-training Settings
The WiFo-M2-LiDAR is trained for $100$ epochs using a batch size of $1024$. The optimization employs the AdamW optimizer with a learning rate of $1\times 10^{-4}$ and a weight decay of $5\times 10^{-4}$. A cosine annealing learning rate scheduler with a warm-up phase is utilized: the learning rate linearly increases from zero to the target learning rate over the first $10$ epochs, then decays following a cosine schedule to zero by the final epoch. The contrastive learning objective is optimized with an initial temperature parameter $\tau_{\rm e}=0.1$ and the output feature dimension $d$ set to $512$. We set the loss weights in Eq. (IV-C2) to $\lambda_{1}=1$, $\lambda_{2}=0.25$, and $\lambda_{3}=0.25$. The WiFo-M2-Img shares the same hyperparameter settings as WiFo-M2-LiDAR, except that the learning rate is set to $5\times 10^{-4}$.
@@ V-C Performance Evaluation
The testing dataset introduced in Section V-A is further divided into a training set and a validation set with a ratio of $4:1$, referred to as the T-T dataset and the T-V dataset, respectively. Note that in all subsequent experiments, the existing schemes are trained on the T-T dataset and their performance is evaluated on the T-V dataset. During this process, the weight of the WiFo-M2 model remains frozen, and only the existing transceiver design networks and the adapter module are trained.
V-C1 Beam Prediction
  • $\bullet$
    Problem Description: The BP task aims to directly predict the optimal beamforming vector $\mathbf{f}^{*}\in\mathcal{F}$ from a predefined codebook $\mathcal{F}=\{\mathbf{f}_{q}\}_{q=1}^{Q}$ at BS using multi-modal sensing. Let $\mathbf{h}\in\mathbb{C}^{1\times N}$ be the CSI of a certain subcarrier. The optimal vector $\mathbf{f}^{*}=\mathop{\arg\max}_{m\in\{1,2,\cdots,Q\}}|\mathbf{h}\mathbf{f}_{m}|^{2}$ results in the maximum effective channel gain and should be chosen for transmission. The BP task can be formulated as: $\mathbf{f}^{*}=\mathcal{Q}_{\rm BP}(\mathbf{M})$, where $\mathcal{Q}_{\rm BP}(\cdot)$ represents a certain BP scheme and $\mathbf{M}$ is the multi‑modal sensory data.
    Unlike the other three transceiver design modules, BP entirely relies on multi-modal sensing. Therefore, the WiFo-M2 model is directly connected to a $2$-layer multilayer perceptron (MLP) network to output the optimal beam index. During fine-tuning on the T-T dataset, only the MLP is trained. Under this setting, the BP performance intuitively reflects the effectiveness of the OOB channel features.
  • $\bullet$

### v2

@@ IV-B2 LiDAR Processing Pipeline
IV-C Network Architecture and Working Flow of WiFo-M2
@@
IV-C1 CSI Feature Extraction
For CSI feature extraction, we employ the pre-trained WiFo model [23] as a powerful space-temporal-frequency (STF) feature extractor. Given an input CSI data $\mathbf{H}_{t}\in\mathbb{C}^{K\times N}$ at timestamp $t$, the WiFo model processes it through its transformer-based encoder. The output is a fixed-dimensional feature vector $\mathbf{z}^{\text{C}}_{t}\in\mathbb{R}^{d}$.
IV-C2 Image and LiDAR Feature Extraction
@@
where $f_{\text{ResNet}}$ and $f_{\text{PointNet}}$ denote the ResNet-34 and PointNet backbone networks respectively. We denote the frame-level features for image and LiDAR within a time window $T-n_{\rm I}\Delta t_{\rm I},T]$ or $T-n_{\rm L}\Delta t_{\rm L},T]$ as $\bm{u}_{t}^{\text{I}}\in\mathbb{R}^{d}$ and $\bm{u}_{t}^{\text{L}}\in\mathbb{R}^{d}$, respectively.
Stage 2: Temporal Feature Modeling. To capture temporal dependencies, both modalities employ the GRU network that learns the underlying dynamics of the environment. The GRU processes the sequences of intermediate frame-level features $\{\bm{u}^{\rm x}_{t}\}_{t=T-n_{\rm x}\Delta t_{\rm x}}^{T},{\rm x}\in\{\rm L,\rm I\}$ to produce a sequence of hidden states, which are then projected to the final feature space. Let $\Delta\tau$ be the PHY action execution interval. The resulting projected feature sequence corresponds to the prediction window $T,T+\Delta\tau,\cdots,T+\Delta t_{\rm x}]$, i.e., from the current sensing timestamp $T$ up to the moment when the next sensing frame is captured. This operation can be expressed as:
@@
Stage 3: Multi-Timestamp Feature Generation and Alignment. A key design of WiFo-M2’s architecture is that the GRU generates channel-aware features for multiple future timestamps. During pre-training, the sequence of final features $\{\bm{v}^{\rm x}_{\tau}\}_{\tau=T}^{T+\Delta t_{\rm x}},{\rm x}\in\{\rm L,\rm I\}$ generated by WiFo-M2 are individually aligned with the CSI feature at the corresponding timestamp. For a given timestamp $\tau$, a certain sensing feature $\bm{v}^{\rm{x}}_{\tau}$ is selected and paired with the CSI feature $\bm{z}^{\text{C}}_{\tau}$ from the same timestamp to form a positive pair for CL (detailed in the next section). As shown in Fig. 5, this design enables WiFo-M2 to learn to generate temporally aligned features at specific future timestamps, which can then be directly compared with CSI features from the corresponding time in CL objective, ensuring precise cross-modal temporal alignment.
IV-D Pre-training Strategy for WiFo-M2: ContraSoM
This section presents the proposed ContraSoM pre-training strategy in detail, covering the modality-specific data-augmentation techniques together with the overall contrastive learning pipeline. By aligning multi-modal environment sensing and wireless channel in the latent space, ContraSoM essentially performs the SoM processing outlined in [4].
IV-D1 Modality-Specific Data Augmentation
@@
The augmentation process begins with a forward noising step. Taking the final LiDAR feature at a specific time step $i$, denoted as $\bm{v}^{\rm L}_{i}$, as an example, we gradually inject Gaussian noise into this feature over a series of $M$ discrete steps. This forward noising process is defined as a Markov chain:
@@
$\tilde{\bm{v}}^{\rm L}_{m}=\sqrt{\bar{\alpha}_{m}}\,\bm{v}^{\rm L}_{i}+\sqrt{1-\bar{\alpha}_{m}}\,\epsilon,\quad\epsilon\sim\mathcal{N}(0,\mathbf{I})$ (11)
@@
For efficient feature generation in the reverse process, we use a deterministic skip-step reverse diffusion sampler. Starting from a noisy feature $\tilde{\bm{v}}^{\rm L}_{m}$, we update with stride $\Delta m=\max(\lfloor m/u\rfloor,1)$ (where $u=8$) as
$\tilde{\bm{v}}^{\rm L}_{m-\Delta m}=$ $\sqrt{\bar{\alpha}_{m-\Delta m}}\left(\frac{\tilde{\bm{v}}^{\rm L}_{m}-\sqrt{1-\bar{\alpha}_{m}}\,\epsilon_{\theta}(\tilde{\bm{v}}^{\rm L}_{m},m)}{\sqrt{\bar{\alpha}_{m}}}\right)$
$+\sqrt{1-\bar{\alpha}_{m-\Delta m}}\,\epsilon_{\theta}(\tilde{\bm{v}}^{\rm L}_{m},m).$ (12)
After $U=24$ reverse steps, we obtain the augmented feature $\tilde{\bm{v}}^{\rm L}_{i}$. To enhance diversity, we generate two distinct augmented views $\tilde{\bm{v}}_{i}^{\rm L(1)}$ and $\tilde{\bm{v}}_{i}^{\rm L(2)}$ by initializing from different noise levels and noise draws, then applying the same reverse update:
$\tilde{\bm{v}}^{\rm L}_{m_{j}}=\sqrt{\bar{\alpha}_{m_{j}}}\,\bm{v}^{\rm L}_{i}+\sqrt{1-\bar{\alpha}_{m_{j}}}\,\epsilon_{j},\quad j\in\{1,2\},$ (13)
where $m_{1},m_{2}\sim\mathcal{U}\{200,400\}$ and $\epsilon_{1},\epsilon_{2}\sim\mathcal{N}(0,\mathbf{I})$. Then, run the above deterministic skip-step reverse updates for $U$ steps to obtain
$\tilde{\bm{v}}_{i}^{\rm L(1)}=\mathcal{G}_{\text{rev}}(\tilde{\bm{v}}^{\rm L}_{m_{1}}),\qquad\tilde{\bm{v}}_{i}^{\rm L(2)}=\mathcal{G}_{\text{rev}}(\tilde{\bm{v}}^{\rm L}_{m_{2}}),$
where $\mathcal{G}_{\text{rev}}(\cdot)$ denotes the $U$-step deterministic reverse sampler defined above. The two independent pairs $(m_{1},\epsilon_{1})$ and $(m_{2},\epsilon_{2})$ yield two complementary augmented views.
IV-D2 Contrastive Learning Objectives in WiFo-M2
@@
For the LiDAR modality, we incorporate both the original LiDAR feature $\bm{v}^{\rm L}$ and the two diffusion-augmented features $\tilde{\bm{v}}^{\rm L(1)}$, $\tilde{\bm{v}}^{\rm L(2)}$ into the CL framework. The overall objective consists of two components: a contrastive loss $\mathcal{L}^{\rm L}_{\text{contra}}$ and a diffusion training loss $\mathcal{L}_{\text{diff}}$. Specifically, we formulate a triplet contrastive loss to enforce consistency between three pairs: the original LiDAR features and CSI features, and the augmented LiDAR features and CSI features:
@@
$\mathcal{L}_{\text{diff}}=\mathbb{E}_{m,\bm{v}^{\rm L},\epsilon}\left[\|\epsilon-\epsilon_{\theta}(\tilde{\bm{v}}^{\rm L}_{m},m)\|^{2}\right],$ (18)
with $m\sim\mathcal{U}\{1,\dots,T\}$ and $\epsilon\sim\mathcal{N}(0,\mathbf{I})$.
@@ V-A Datasets
Pre-training Dataset M3SC [26] Intersection traffic $28$ I1 $32\times 1$ $0.2$ $512$ $7,048$
I2 $64\times 1$ 11 $1024$ $13,888$
I3 $16\times 8$ $0.8$ $512$ $7,088$
I4 $16\times 8$ $0.4$ $512$ $7,168$
I5 $8\times 8$ $0.8$ $1024$ $14,048$
I6 $8\times 4$ 11 $256$ $3,484$
SynthSoM [27] Dense building $4.95$ B1 $16\times 16$ $0.05$ $256$ $9,080$
B2 $16\times 8$ $0.05$ $512$ $16,784$
B3 $8\times 8$ $0.05$ $512$ $14,816$
B4 $8\times 4$ $0.1$ $512$ $18,824$
B5 $4\times 4$ $0.1$ $1024$ $38,064$
@@
We construct a multi-modal pre-training dataset that includes LiDAR, RGB images, and CSI data. The CSI data is collected from the links between the BS and the users. In all datasets, the users are equipped with a single antenna, while the BS employs either a UPA or ULA configuration, with adjacent antenna spacing set to half the wavelength at the central frequency. The data is sourced from the intersection traffic scenario in the M3SC dataset [26] and the dense building scenario in the SynthSoM dataset [27]. Various communication system configurations are considered, covering a wide range of subcarrier numbers, bandwidths, and antenna sizes. The detailed simulation configurations of the pre-training datasets are presented in Table I.
We also construct a testing set, with detailed simulation configurations shown in Table II. The testing set consists of two parts: $7$ datasets for seen scenarios and $3$ datasets for unseen scenarios. First, to validate the in-distribution performance of WiFo-M2, we construct datasets L1 to L7, which are selected from the same links as the pre-training datasets. Among these, datasets L1, L5, and L6 are formed by merging several links with the same number of antennas and subcarriers but smaller data volumes. To test the cross-scenario generalization ability of WiFo-M2, we construct datasets C1, S1, and V1 based on three new scenarios not included in the pre-training datasets, namely, the campus scenario in the SynthSoM-Twin dataset [29], the “dist_cam” scenario in the ViWi dataset [30], and the Scenario 31 in the DeepSense-6G dataset [31], respectively. Notably, these three datasets operate at entirely new carrier frequencies, and their images and LiDAR point clouds exhibit visual appearances and spatial layouts unseen in the pre-training data. This pronounced gap between training and evaluation domains highlights the strong universality and generalization capability of WiFo‑M2. Fig. 6 provides an overview of the multi-modal environment sensory data examples from all the aforementioned scenarios.
@@ V-B Network and Pre-training Settings
The WiFo-M2-LiDAR is trained for $100$ epochs using a batch size of $1024$. The optimization employs the AdamW optimizer with a learning rate of $1\times 10^{-4}$ and a weight decay of $5\times 10^{-4}$. A cosine annealing learning rate scheduler with a warm-up phase is utilized: the learning rate linearly increases from zero to the target learning rate over the first $10$ epochs, then decays following a cosine schedule to zero by the final epoch. The contrastive learning objective is optimized with an initial temperature parameter $\tau_{\rm e}=0.1$ and the output feature dimension $d$ set to $512$. We set the loss weights in Eq. (IV-D2) to $\lambda_{1}=1$, $\lambda_{2}=0.25$, and $\lambda_{3}=0.25$. The WiFo-M2-Img shares the same hyperparameter settings as WiFo-M2-LiDAR, except that the learning rate is set to $5\times 10^{-4}$.
@@ V-C Performance Evaluation
In this work, we select four representative PHY actions: beam prediction, pilot-based channel estimation, channel interpolation and channel prediction, which together span the whole CSI pipeline from acquisition to utilization and underpin reliable network connectivity, providing a comprehensive test bed for our framework.
The testing dataset is further divided into a training set and a validation set with a ratio of $4:1$, referred to as the T-T dataset and the T-V dataset, respectively. Note that in all subsequent experiments, the existing schemes are trained on T-T dataset and their performance is evaluated on T-V dataset. During this process, the weight of the WiFo-M2 remains frozen, and only the existing schemes and the adapter module are trained.
V-C1 PHY Action 1-Beam Prediction
  • Problem Description: BP refers to directly predicting the optimal beamforming vector $\mathbf{f}^{*}\in\mathcal{F}$ from a pre-defined codebook $\mathcal{F}=\{\mathbf{f}_{q}\}_{q=1}^{Q}$ using multi-modal sensing. Let $\mathbf{h}\in\mathbb{C}^{1\times N}$ be the CSI of a certain subcarrier. The optimal vector $\mathbf{f}^{*}=\mathop{\arg\max}_{m\in\{1,2,\cdots,Q\}}|\mathbf{h}\mathbf{f}_{m}|^{2}$ results in the maximum effective channel gain and should be chosen for transmission. The BP action can be formulated as: $\mathbf{f}^{*}=\mathcal{Q}_{\rm BP}(\mathbf{M})$, where $\mathcal{Q}_{\rm BP}(\cdot)$ represents a certain BP scheme and $\mathbf{M}$ is the multi‑modal sensory data.
    Unlike the other three PHY actions, BP entirely relies on multi-modal sensing. Therefore, the WiFo-M2 model is directly connected to a $2$-layer multilayer perceptron (MLP) network to output the optimal beam index. During fine-tuning on the T-T dataset, only the MLP is trained. Under this setting, the BP performance intuitively reflects the effectiveness of the OOB channel-aware features.